• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 16:00:51 on March29,2005

    Entry number 142723

    keyword=Shift Summary


    Run 3033

    At 8:30 stopped taking data for MCC beam studies.

    Set prescalers to take cosmic data.
    (set the T6 and T7 prescalers to 1)

    3034 Junk Run (needed to reset TS)
    3035 Cosmics
    3037 Cosmics
    3038 Cosmics
    3039 Cosmics
    3040 Cosmics
    3041 Cosmics

    At 11:45 the beam was returned and we did a visual check using the
    BeO target. After convincing ourself that the beam was not hitting the frame, we moved to the 12C target and continued taking (e,e'pN)
    data.

    Carbon Production Runs

    3042 Production 12:10 13:00
    3043 Production 13:00 14:01
    3044 Production 14:01 14:48
    3045 Production 14:48 15:34
    3046 Production 15:34 (through end of shift)